# Good Noise Makes Good Edits: A Training-Free Diffusion-Based Video Editing with Image and Text Prompts

### v1

@@ 2.3.2 $\mathbf{\rho}$-start sampling
Preserving consistency with the source video is critical in video editing, particularly for maintaining visual structure and motion save ; camel ; deco . We hypothesize that initiating the sampling process from a less perturbed latent, closer to the source video, facilitates the preservation of these characteristics.
To this end, we propose an alternative sampling strategy that modifies the standard DDIM schedule by starting the denoising process from a less noisy latent representation, thereby improving structural and temporal fidelity. In the conventional reduced DDIM schedule ddim , sampling proceeds along a trajectory ${{\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\nu}}}$, starting from a highly noisy latent ${\bm{z}}_{\tau_{\nu}}$. In contrast, our approach begins from an earlier timestep, ${\bm{z}}_{\tau_{\rho}}$, where $\rho<\nu$. This shortens the denoising trajectory to only $\rho$ steps, $\{{\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\rho}}\}$, facilitating the retention of more information from the source video.
@@ 2.4 Zero image guidance
To enhance color fidelity, we introduce zero image guidance, helping better align with the appearance of the reference image. Our refined guidance strategy blends the conventional negative image embeddings CFG ; glide , represented by a zero vector of the same size as the CLIP image embedding, with a zero image guidance, which refers to the CLIP image embedding derived from an all-zero image. The final negative image guidance $\bar{{\bm{c}}}_{\text{img}}$ is computed as a weighted sum of the two, controlled by a blending factor $\gamma$, and can be formally expressed as:
$\bar{{\bm{c}}}_{\text{img}}=\gamma\cdot E_{I}({\mathbf{I}}_{0})+(1-\gamma)\cdot\mathbf{0},$ (8)
where $\gamma\in[0,1]$ is a hyperparameter, $\mathbf{0}\in\mathbb{R}^{d}$ is a zero vector with the same size as the CLIP image embedding, $E_{I}(\cdot)$ represents the CLIP image encoder, and ${\mathbf{I}}_{0}\in\mathbb{R}^{C\times H\times W}$ is an all-zero image input. This formulation enables a tunable trade-off between the two types of negative guidance, allowing finer control over the color intensity of the generated object. The resulting $\bar{{\bm{c}}}_{\text{img}}$, together with the conventional $\bar{{\bm{c}}}_{\text{txt}}$, serves as negative guidance during the sampling process in Eq. 3.
@@ 3 Experiments
3.1 Experimental setup
Evaluation metrics. We conduct evaluations on 18 pairs of video and reference image, using videos collected from DAVIS dataset davis and the dataset released by Video-P2P videop2p . To assess how well the generated videos align with the given text and image prompt, we use CLIP-T score and DINO score. Specifically, CLIP-T score measures the alignment between the target text prompt and the generated videos using CLIP text encoder clip , while DINO score calculates the similarity between the reference image and the generated video using DINO dino . To evaluate temporal consistency, we extract CLIP image embeddings for each frame in the edited video using the CLIP image encoder clip and measure the cosine similarity between embeddings of adjacent frames.
Baseline models. We compare our proposed method with two baselines: Video-P2P videop2p and Make-A-Protagonist (Make-A-Pro) make_a_pro . To the best of our knowledge, Make-A-Pro is the only baseline that directly aligns with our problem setting and offers publicly released code among existing methods. However, in order to comprehensively demonstrate the effectiveness of our image prompt approach, we also include Video-P2P, a state-of-the-art text-driven video editing model, as an additional baseline.
Implementation details. We use AnimateDiff animatediff as the base T2V model, IP-Adapter ipadapter v1.5 for the reference image encoding, and SAM2 sam2 for obtaining masks. We conduct experiments with videos of various resolutions (320$\times$576, 480$\times$480, 512$\times$512, 600$\times$600 and 640$\times$640) and different numbers of frames (8 and 16). For text prompts, we use the same descriptions when comparing our method with Video-P2P videop2p . In contrast, for Make-A-Pro make_a_pro , we follow their strategy by extracting protagonist information using BLIP-2 blip2 and generating prompts via visual question answering. Additionally, since Video-P2P videop2p only supports square outputs, we paste its generated video back into the corresponding regions of the source video to enable a comparison. All experiments are conducted on a single NVIDIA A100 GPU with 40GB memory.
3.2 Experimental results
3.2.1 Comparisons with state-of-the-art methods
Quantitative and qualitative results. We quantitatively evaluate our method with a focus on alignment with text and image prompts, visual feature similarity, and temporal consistency across video frames. As shown in Table 1, our method consistently outperforms all baselines across all evaluation metrics, both automatic and human. In terms of model scoring metrics, our method achieves the highest CLIP-T, DINO, and Temp scores, indicating superior text alignment, object similarity, and temporal consistency, respectively. These trends are further supported by human evaluation. Our method receives the lowest scores in Image, Text, and Temp categories, demonstrating that it generates objects most similar to the target, aligns most faithfully with the input text prompts, and maintains the most consistent appearance across frames. In addition, our method is rated highest in Overall quality, clearly confirming its effectiveness and perceptual superiority over the baselines.
Qualitative comparisons further demonstrate the effectiveness of our method in accurately preserving both motion and identity across diverse editing scenarios, outperforming the baselines. In the left example of Fig. 6, Make-A-Pro make_a_pro preserves the motion from the source video but fails to accurately reflect the identity of the reference image, even struggling to edit the dog into a fox. Video-P2P videop2p partially captures the identity of the fox but fails to preserve motion, resulting in an incorrect number of objects, with two foxes appearing in a single frame. In contrast, our method produces the most accurate edited video, successfully maintaining both the motion from the source video and the identity of the reference object. The right example in Fig. 6 features a source video with less dynamic movement. In this case, all methods preserve motion, but only our method correctly generates a white duck as specified in both the text condition and reference image. This example highlights the ability of our method to effectively handle both image and text prompts in video editing tasks.
Table 1: Quantitative comparison results. Quantitative metrics include CLIP-Text similarity (CLIP-T), visual feature similarity (DINO), and temporal consistency (Temp). Human evaluations measure the text-video alignment (Text), visual fidelity to the target object (Image), motion consistency with the source video (Motion), temporal consistency (Temp), and overall visual quality (Overall).
Method Model scoring metrics Human evaluations
CLIP-T $\uparrow$ DINO $\uparrow$ Temp $\uparrow$ Text $\downarrow$ Image $\downarrow$ Motion $\downarrow$ Temp $\downarrow$ Overall $\downarrow$
Video-p2p videop2p 0.300 0.393 0.961 2.01 2.05 2.27 1.99 2.17
Make-A-Pro make_a_pro 0.294 0.362 0.958 2.66 2.61 1.99 2.64 2.57
ImEdit (Ours) 0.326 0.474 0.966 1.33 1.34 1.74 1.37 1.26
Figure 6: Qualitative comparison results on DAVIS dataset. We compare the generation quality of our method against baseline methods. While the baseline methods fail to incorporate key visual features from the reference image, our method effectively transfers the target object into the source video, faithfully preserving its appearance across frames.
User study. We conduct a user study based on five evaluation criteria: (1) image alignment, which measures how well the object in the generated video resembles the one in the reference image; (2) text alignment, which assesses how accurately the generated video matches the given text condition; (3) motion alignment, which evaluates how well the motion in the generated video aligns with that of the source video; (4) temporal consistency, which examines the consistency of the object and background across frames; and (5) overall quality, which considers all the aforementioned factors to determine how natural and realistic the generated video appears. We ask 32 participants to evaluate 9 comparisons each, ranking the given results based on these criteria. We compute the average rankings, and Table 1 shows that our method ranks highest across all criteria.
3.2.2 Ablation study
Quantitative results. We quantitatively evaluate the contribution of each component in our method, as shown in Table 2. Starting from the base model (a), the integration of zero image guidance (b) leads to an overall performance improvement. Incorporating $\rho$-start sampling (c) brings a slight increase in CLIP-Text alignment, while the DINO and Temporal scores slightly decrease, indicating that adjusting the starting sampling point alone is insufficient to resolve the trade-off between motion preservation and appearance fidelity. The inclusion of dual masking (d) contributes to a more substantial gain across all metrics. Finally, our proposed method with dilation (e) achieves the best performance, with notable improvements. These results highlight the effectiveness of the complete method.
Effectiveness of dilated dual masking. We ablate our proposed dilated dual masking by progressively incorporating each masking schemes, as shown in Fig. 7. We begin with the simplest configuration: a single stage editing setup that relies solely on the source mask. When using the source mask only, the generated object is often forced to conform to the contours of the source object, leading to unrealistic shapes, limited fidelity, and artifacts. This issue becomes particularly pronounced when there is a significant shape difference between the source and target objects. As shown in Fig. 7(a), the source object (grey car) is structurally different from the target object; it is shorter in length, taller in height, and more angular than the target object, but these differences are not considered.
Adding an additional stage to generate the rough mask and using only this mask alleviates the problem of the imprecise reflection of the target object’s features. However, Fig. 7(b) reveals that it fails to completely remove the source, as the top part of the grey car still remains. To further address this, we take the union of the source and rough mask. While this reduces residuals from the source object, artifacts persist along the boundaries. Furthermore, fine-grained details of the generated target object are lost due to the strict constraints imposed by the merged mask region, as illustrated in Fig. 7(c).
Our proposed dilated dual masking effectively addresses the aforementioned issues by applying dilation to the masks and this is evident in Fig. 7(d). It allows the generated target object to extend beyond the strict union region, enabling natural shape adaptation and smoother blending into the surrounding context. As a result, the proposed dilated dual masking achieves both structural integrity and visual realism, demonstrating its superiority.
@@ 3.2.2 Ablation study
Table 2: Comprehensive ablation study of ImEdit. We evaluate the contribution of each component to the overall performance of our method. (a) Base model. (b) Zero image guidance. (c) $\rho$-start sampling. (d) Dual masking. (e) Dilation.
Method CLIP-Text $\uparrow$ DINO $\uparrow$ Temp $\uparrow$
(a) 0.302 0.361 0.958
(a) + (b) 0.304 0.405 0.960
(a) + (b) + (c) 0.309 0.401 0.955
(a) + (b) + (c) + (d) 0.316 0.435 0.958
(a) + (b) + (c) + (d) + (e) (Ours) 0.326 0.474 0.966
Figure 7: Ablation study on the proposed dilated dual masking. (a) shows the output of the first stage using the source video $\mathcal{V}_{s}$ and source object mask, while (b)-(d) present outputs of the second stage from the intermediate video $\mathcal{V}_{r}$ using different masking strategies. Masks are overlaid on the corresponding video. The proposed dilated dual masking (d) achieves the most accurate edits, underscoring its effectiveness for the object replacement.
Figure 8: Effect of varying the $\rho$ value. The rightmost column shows the final output after the second stage, while all other columns, excluding the reference image and source video, show the results of the first stage under different values of $\rho$. By adjusting sampling start point, we balance how strongly the appearance of the target object is retained versus the source motion. Lower $\rho$ emphasizes the owl’s visual traits, while higher $\rho$  increasingly preserves the parrot’s original movement and posture.
@@
Effectiveness of $\rho$-start sampling. As shown in Fig. 8, varying the parameter $\rho$, which determines the sampling start point, enables control over the balance between the target appearance and source motion. Using a larger $\rho$ leads the model to fails to accurately follow the motion and spatial layout of the source video. For example, while the parrot in the source video is facing left, the owl in the edited video is facing forward or at an incorrect angle. In contrast, using a smaller $\rho$ allows the model to better capture the motion and spatial dynamics of the source video, resulting in the owl correctly facing sideways, consistent with the source object. However, if $\rho$ is too small, the transformation may become too weak, leading to incomplete replacement of the source object. With an appropriate choice of $\rho$  the edited video successfully balances appearance fidelity and motion preservation, highlighting the importance of temporal alignment in video editing tasks.
@@
Effectiveness of the zero image guidance. We analyze the impact of $\gamma$ in the proposed zero image guidance. As shown in Fig. 9, increasing $\gamma$, which scales the contribution of the zero image guidance (see Eq. (8)), leads to more saturated colors. This suggests that mixing the zero image guidance into the negative image prompt effectively enhances color intensity, providing a simple yet powerful control for visual intensity.
Figure 9: Effect of varying the $\gamma$ value. Increasing $\gamma$ enhances the influence of the zero image embedding, leading to more saturated color reproduction in the edited video.
@@ 4 Conclusion
In this work, we propose ImEdit, a novel zero-shot, training-free video editing method guided by both image and text prompts. ImEdit enables accurate object replacement with strong temporal coherence and achieves superior performance in both visual quality and motion consistency, outperforming existing state-of-the-art approaches with high efficiency and scalability.
@@
This supplementary material complements the main manuscript, providing additional experimental results and implementation details.
  • Appendix A provides an analysis of the effect of the DDIM inversion trajectory length.
  • Appendix B provides an additional quantitative comparison with baseline methods.
@@
  • Appendix D addresses the limitations of our method.
  • Appendix E provides additional ablation results on dilated dual masking.
  • Appendix F provides additional implementation details and user study details.
  • Appendix G presents the overall algorithm of the proposed method.
  • Appendix H provides additional qualitative results.
Appendix A Effect of the DDIM inversion trajectory length
In our latent blending process, the number of DDIM inversion sampling steps is treated as a hyperparameter with two discrete settings, 999 or 50, depending on the desired outcome. Using 999 steps often results in less vivid colors and a relatively noisy background. However, the edited object tends to exhibit more natural shapes and the overall video appears more realistic. In contrast, using 50 steps yields cleaner, more vivid appearances for both the object and background, and significantly reduces DDIM inversion time. Nonetheless, the shape of the edited object may appear less natural.
As shown in Fig. 10, the result with 999 steps shows a polar bear with reduced color saturation and increased background noise while the head and front paws are generated with a more natural structure. Conversely, the result with 50 steps shows a polar bear with more vivid white appearance and cleaner background, but the shape of the head and paws is less natural.
Figure 10: Effect of the reduced inversion trajectory length. We evaluate how varying $\nu$, the number of DDIM steps used in inversion, affects editing quality.
Appendix B Additional quantitative results
Table 3 presents an additional quantitative comparison between ImEdit and existing baseline methods using CLIP-T, DINO, Temp, and PSNR, which respectively assess text-image alignment, the similarity between the reference image and the edited video, temporal consistency, and background preservation. To facilitate a more comprehensive evaluation, we extend the comparison in Table 1 by including FateZero fatezero and a naive variant of our approach, AnimateDiff animatediff combined with IP-Adapter ipadapter , and by adding PSNR as an additional metric. Specifically, AnimateDiff + IP-Adapter refers to the direct combining of our backbone model AnimateDiff with IP-Adapter, without any of the proposed components, as illustrated in Fig. 2. FateZero fatezero is included as a baseline due to its strong editing performance among text-driven video editing methods that operate in a training-free setting, consistent with the setting of our method.
Table 3: Additional quantitative comparison with baseline models. We compare our method (ImEdit) against existing baselines across four metrics: CLIP-Text similarity (CLIP-T), visual feature similarity (DINO), temporal consistency (Temp), and background preservation (PSNR).

### v2

@@ 2.3.2 $\mathbf{\rho}$-start sampling
Preserving consistency with the source video is critical in video editing, particularly for maintaining structural integrity and coherent motion [save, camel, deco]. Inspired by SDEdit [sdedit], we hypothesize that initiating the sampling process from a less perturbed latent, closer to the source video, helps retain both spatial structure and temporal dynamics.
To this end, we propose an alternative sampling strategy that modifies the standard DDIM schedule by starting the denoising process from a less noisy latent representation, thereby improving structural and temporal fidelity (Fig. 4(d)). In the conventional reduced DDIM schedule [ddim], sampling proceeds along a trajectory ${\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\nu}}$}, starting from a highly noisy latent ${\bm{z}}_{\tau_{\nu}}$. In contrast, our approach begins from an earlier timestep, ${\bm{z}}_{\tau_{\rho}}$, where $\rho<\nu$. This shortens the denoising trajectory to only $\rho$ steps, $\{{\bm{z}}_{\tau_{1}},\ldots,{\bm{z}}_{\tau_{\rho}}\}$, better preserving source content overall.
@@
Table 1: Quantitative comparison. We compare VINO against four state-of-the-art models and two naive adaptations of image editing and video generation models for video editing. We measure performance using both automatic metrics and human evaluations to assess text, image, and motion alignment, temporal consistency, and overall quality.
Method Model support Model scoring metrics Human evaluations
Text Image Train-free DINO $\uparrow$ CLIP-T $\uparrow$ Temp $\uparrow$ PSNR $\uparrow$ Image $\downarrow$ Text $\downarrow$ Motion $\downarrow$ Temp $\downarrow$ Overall $\downarrow$
SDEdit-Video [sdedit] 0.431 0.303 0.959 16.3 4.89 5.08 5.93 5.02 5.36
FateZero [fatezero] 0.438 0.311 0.967 20.1 4.59 4.53 4.44 5.06 4.75
Video-P2P [videop2p] 0.424 0.305 0.971 19.0 4.00 4.07 3.68 4.09 3.70
TokenFlow [tokenflow] 0.480 0.325 0.975 19.6 3.78 3.09 2.61 2.50 2.95
AnimateDiff [animatediff]+IPA [ipadapter] 0.399 0.308 0.955 24.7 4.59 5.01 5.24 4.30 4.93
Make-A-Pro [make_a_pro] 0.444 0.309 0.965 16.9 3.93 4.24 3.64 5.00 4.35
[…]
Figure 6: Qualitative comparison. We compare the generation quality of VINO against four state-of-the-art baseline models. VINO better preserves reference appearance across frames, while baselines struggle to retain key visual features of the target object.
@@ 2.4 Zero image guidance
To enhance color fidelity, we present zero image guidance (Fig. 4(c)). It blends the conventional negative image embeddings [CFG, glide], defined by a zero vector of the same size as the CLIP image embedding, with a zero image guidance, which refers to the CLIP image embedding derived from an all-zero image. The final negative image guidance $\bar{{\bm{c}}}_{\text{img}}$ is computed as a weighted sum of the two, controlled by a blending factor $\gamma$:
$\bar{{\bm{c}}}_{\text{img}}=\gamma\cdot\text{CLIP}_{\text{img}}({\mathbf{I}}_{0})+(1-\gamma)\cdot\mathbf{0},$ (11)
where $\gamma\in[0,1]$ is a hyperparameter, $\mathbf{0}\in\mathbb{R}^{d}$ is a zero vector with the same size as the CLIP image embedding, $\text{CLIP}_{\text{img}}(\cdot)$ is the CLIP image encoder, and ${\mathbf{I}}_{0}\in\mathbb{R}^{C\times H\times W}$ is an all-zero image. This formulation offers a parameterized trade-off between the two types of negative guidance, allowing finer control over the color intensity of the generated object. The resulting $\bar{{\bm{c}}}_{\text{img}}$, combined with the conventional $\bar{{\bm{c}}}_{\text{txt}}$, is used as negative guidance during the sampling process in Eq. 6.
Figure 7: Ablation study on the dilated dual masking. The output of the first stage using the source video $\mathcal{V}_{s}$ and source object mask is shown in (a). Subsequently, the outputs of the second stage from the intermediate video $\mathcal{V}_{r}$ using different masking strategies are presented in (b)-(d). Masks are overlaid on the corresponding video. The proposed dilated dual masking (d) achieves the most accurate edits.
Figure 8: Ablation study on the effect of varying the $\rho$ value.
@@ 3 Experiments
Dataset. We evaluate VINO on 20 video-image pairs. Videos are sourced from DAVIS [davis], Video-P2P [videop2p], and a curated set covering various object categories (e.g., humans, animals, everyday items) to ensure broad semantic and visual diversity. Reference images include both real and synthetic samples generated by models such as GPT-4o and Gemini.
Baseline models. We compare VINO against six representative video editing methods, as summarized in Table 1, which use text or image prompts for semantic appearance control. SDEdit-Video [^2] We use the HuggingFace implementation of SDEdit extended to video. [sdedit], FateZero [fatezero], Video-P2P [videop2p], and TokenFlow [tokenflow] represent text-guided video editing methods. To our knowledge, Make-A-Protagonist (Make-A-Pro) [make_a_pro] is the only publicly available method that supports both image and text conditioning, making it the most directly comparable to VINO. AnimateDiff [animatediff] +IP-Adapter (IPA) [ipadapter] serves as a modular baseline that integrates a pretrained T2V model with an image-guided adapter, illustrating the limitations of naive fusion strategies.
Implementation details. Our main experiments adopt AnimateDiff [animatediff] as the base T2V model, IP-Adapter v1.5 [ipadapter] for reference image conditioning, and SAM2 [sam2] for mask generation. All experiments are conducted on a single NVIDIA A100 GPU. Additional implementation details, including hyperparameters, baseline setups, and extensions to other T2V backbones, are provided in Appendix E.1.
Due to space constraints, we present representative qualitative comparisons and ablation results in the main paper, with full qualitative results provided in Appendix B.
Figure 9: Ablation study on the image guidance scale $\gamma$.
3.1 Comparisons with state-of-the-art methods
Quantitative results. We evaluate VINO using four metrics: DINO score measures visual alignment via cosine similarity between features of the reference image and generated video [dino]. CLIP-Text score quantifies text alignment by computing cosine similarity between CLIP embeddings of the generated video and the text prompt [clip]. Temporal score assesses motion consistency through cosine similarity between CLIP image embeddings of adjacent frames [clip]. PSNR score measures how accurately the background of the source video is preserved by computing pixel-level similarity between the generated and source frames.
VINO achieves the highest DINO score among all baselines, indicating strong visual fidelity to the reference appearance. In CLIP-Text and temporal scores, it performs comparably to TokenFlow while outperforming other methods. Notably, VINO is compatible with existing methods such as TokenFlow and can serve as an additional appearance control module. This composition yields further performance gains, as detailed in the Appendix A. For PSNR, AnimateDiff+IPA and VINO noticeably outperform the other methods. The small gap between them, despite both using the same latent blending strategy, arises from the edited area. AnimateDiff+IPA modifies a comparatively smaller region, whereas VINO employs dilated dual masking, which introduces a larger generation area, leading to a marginally lower PSNR.
User study. We conduct a user study with 25 participants to assess perceptual quality across five criteria: image alignment, text alignment, motion alignment, temporal consistency, and overall quality. VINO consistently outperforms all baselines across all criteria, demonstrating strong perceptual fidelity in both appearance and motion.
Qualitative results. Qualitative comparisons, shown in Fig. 6, further demonstrate the effectiveness of VINO in preserving object appearance and motion across various editing scenarios, consistently outperforming baselines.
In the man-to-panda translation, baselines either fail to generate a panda or produce artifacts or distortions, while VINO yields a clean and consistent result. A similar trend appears in the duck-to-chicken translation, where most methods generate generic chicken-like color rather than matching the reference, whereas VINO preserves its specific visual traits. In the rabbit-to-fox translation, most baselines struggle to maintain appearance fidelity or temporal coherence. Make-A-Pro, in particular, suffers from temporal flickering, including disappearing or reappearing limbs. In contrast, VINO maintains the correct fox appearance while preserving the background, producing coherent structure and stable motion throughout the video.
@@ 3.1 Comparisons with state-of-the-art methods
Table 2: Comprehensive ablation study of VINO. We evaluate the contribution of each component to the overall performance of our method: (a) base model, (b) $\rho$-start sampling, (c) dual masking, (d) dilation, and (e) zero image guidance.
Method DINO $\uparrow$ CLIP-T $\uparrow$ Temp $\uparrow$
(a) 0.399 0.308 0.955
(a) + (b) 0.429 0.311 0.953
(a) + (b) + (c) 0.466 0.316 0.954
(a) + (b) + (c) + (d) 0.474 0.316 0.955
[…]
3.2 Ablation study
Effectiveness of dilated dual masking. We conduct an ablation study on our dilated dual masking by progressively applying different masking schemes (Fig. 7). Using only the source mask constrains the generation to the contours of the source object, often resulting in unrealistic shapes and artifacts when the source and target differ significantly (Fig. 7(a)). Applying only the rough mask alleviates this issue but leaves visible remnants of the source object (Fig. 7(b)). Taking the union of both masks suppresses these remnants but introduces boundary artifacts and limits the detail due to the rigid mask constraint (Fig. 7(c)).
In contrast, our dilated dual masking effectively mitigates these limitations by dilating both masks before fusion, enabling the generated object to extend beyond the strict union region. This flexibility promotes more natural shape adaptation and smoother blending with the surrounding context (Fig. 7(d)). As a result, our dilated dual masking achieves both structural coherence and visual realism.
@@ 3.2 Ablation study
Effectiveness of $\rho$-start sampling. As shown in Fig. 8, adjusting $\rho$, which controls the sampling start point, enables a trade-off between target appearance and source motion. Larger $\rho$ values cause the model to deviate from the source motion and layout. For example, the parrot facing left in the source video becomes an owl facing forward in the edited result. Smaller $\rho$ values better preserve the source motion and spatial dynamics, producing an owl that correctly faces sideways. However, overly small $\rho$ weakens the transformation, leading to incomplete replacement of the source object. An appropriate choice of $\rho$ achieves a balanced edit, maintaining both appearance fidelity and motion preservation.
@@
Effectiveness of zero image guidance. We analyze the impact of $\gamma$ in our zero image guidance. As shown in Fig. 9, increasing $\gamma$, which scales the strength of the zero image guidance (see Eq. (11)), leads to more saturated colors. This suggests that incorporating the zero image guidance into the negative image effectively enhances color intensity, providing a simple yet powerful control for visual appearance.
Quantitative results. Table 2 summarizes the effect of each component in our method. As we progressively add components (a)–(e), DINO and CLIP-Text scores increase monotonically, indicating that the proposed modules jointly improve appearance transfer from the reference image as well as adherence to the text prompt. In contrast, the temporal score remains relatively stable for variants (a)–(d), but improves significantly once zero image guidance (e) is introduced, leading our complete model to achieve the best performance across all metrics. These results quantitatively demonstrate that each component contributes effectively to temporally coherent video editing conditioned on both image and text.
@@ 4 Conclusion
In this work, we propose VINO, a novel zero-shot, training-free video editing method guided by both image and text prompts. It enables accurate object replacement with strong temporal coherence, outperforming state-of-the-art methods in visual quality and motion consistency, and efficiency.
@@
This supplementary material complements the main manuscript, providing additional experimental results and implementation details that could not be included because of space limitations.
  • Appendix A illustrates the applicability of our method across various text-to-video models.
  • Appendix B includes complete qualitative comparison and ablation studies.
@@
  • Appendix D discusses the limitations of our method.
  • Appendix E details additional experimental settings.
  • Appendix F presents the overall algorithm of our method.
  • Appendix G showcases additional qualitative results.
Appendix A Applicability to various T2V models
As shown in Table 3, our method is compatible with various text-to-video (T2V) models, enabling training-free video editing conditioned on both images and text. We evaluate the compatibility of our approach by extending the text-based generative model AnimateDiff [animatediff] and the text-based editing model TokenFlow [tokenflow] with our proposed components, including dilated dual masking, $\rho$-start sampling, zero-image guidance, and the image-guided module.
Integrating our method into AnimateDiff yields higher scores across most metrics, demonstrating substantial improvements in structural consistency and visual similarity to the reference image. When applied to TokenFlow, DINO and PSNR also improve, indicating the integration strengthens reference alignment even in the text-based editing models.
However, our method shows a lower temporal consistency score than that of text-prompt-only baselines. This reflects a trade-off inherent to image-based conditioning, where aligning each frame to a high-fidelity reference image introduces larger pixel-level variations, even though the edits remain perceptually coherent. Despite this trade-off, the consistent gains achieved with only marginal drops in temporal consistency across both generation and editing models underscore the applicability and generality of our method.
Figure 10: Full qualitative comparison with baselines. We provide an extended version of Fig. 6 from the main manuscript by including results from SDEdit-Video [sdedit] and AnimateDiff [animatediff]+IP-Adapter [ipadapter].
Figure 11: Full qualitative results of the ablation study on dilated dual masking. Additional frames corresponding to Fig. 7 in the main manuscript are provided to further illustrate the consistent transformation of the object across time.
Appendix B Full qualitative comparison and ablations
This section presents the full qualitative comparisons with baseline methods and qualitative ablation results that complement those in the main paper. Appendix B.1 includes complete qualitative comparison against baseline approaches, Appendix B.2 provides full results for ablation studies.
Table 3: Quantitative evaluation of applicability to various T2V models. We integrate our proposed components into AnimateDiff and TokenFlow. The integration improves reference alignment and image quality with minimal trade-offs across different architectures.